OpenStudy (anonymous):

Will fan and medal Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

X - 2y = 14 X + 3y = 9

OpenStudy (anonymous):

are you there??

OpenStudy (anonymous):

yes but don't get it

OpenStudy (anonymous):

(multiply the second equation by -1) X - 2y = 14 -1( x+ 3y = 9)

OpenStudy (anonymous):

Now add them up, now X cancels and you have Y - 2y - 3y = 14 - 9 -5y = 5 ==> y= -1

OpenStudy (anonymous):

Now add them up, now X cancels and you have Y - 2y - 3y = 14 - 9 -5y = 5 ==> y= -1

OpenStudy (anonymous):

do you get it ???

OpenStudy (anonymous):

so b would the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well your welcome see you another time